OpenStudy (anonymous):

Choose the equation of the line passing through the point (-3, -1) and perpendicular to y =1 over 2 x + 3. - I picked y = -2x +5?

OpenStudy (anonymous):

y= (2/3)x+1

OpenStudy (anonymous):

y = -2x - 7 y = -2x +5 y = 2x + 5 y = 2x - 7 these are the answers they give you